OpenStudy (anonymous):

You give your 30 students a pre-test for their knowledge of statistics on the first day of class for the semester. Their mean score was 40% correct (s = 5.00). Then on the comprehensive final exam you note that their mean score was 70% correct (s = 6.00). The standard deviation of the difference between the 2 means (40 vs. 70) is 4.00. What is the confidence interval for the mean difference in statistics knowledge before and after the course? What is the effect size in comparing knowledge students had before and after the course?

OpenStudy (perl):

do you have the formula

OpenStudy (anonymous):

Yes. Hold on.

OpenStudy (anonymous):

CI 0.95 = XD + t 0.05 (sX ) = XD + t 0.05 (s / SR [ N ] ) This is the formula for the confidence interval.

OpenStudy (perl):

what is XD

OpenStudy (anonymous):

X Bar and I forgot what the number is called when it's little and beneath the Big number? Sorry.

OpenStudy (anonymous):



OpenStudy (anonymous):

Looks like that.

OpenStudy (perl):

ok x bar d stands for the difference of the means

OpenStudy (perl):

sample difference mean

OpenStudy (anonymous):

Okay. I see.

OpenStudy (perl):

and you need to find the t score for .05

OpenStudy (perl):

for 29 degree of freedom the t score is 2.0452

OpenStudy (perl):

so this gives us CI 0.95 = XD + t 0.05 (sX ) = XD + t 0.05 (s / SR [ N ] ) = (70-40) + 2.0452 * 4 / sqrt(30)

OpenStudy (perl):

actually it is an interval ((70-40) - 2.0452 * 4 / sqrt(30) , (70-40) + 2.0452 * 4 / sqrt(30)

OpenStudy (anonymous):

I was using the s= 5.00 and the s= 6.00 in my equation. haha

OpenStudy (perl):

(28.506, 31.494)

OpenStudy (perl):

the next question, can you double check for typo, i dont understand the english

OpenStudy (perl):

What is the effect size in comparing knowledge students had before and after the course?

OpenStudy (anonymous):

Yeah. That does sound confusing. I think it might be asking to use Cohen's equation.